OpenStudy (anonymous):

Suzie is thinking of 2 positive numbers. One of them is 4 more than the other, and the sum of theirs squares is 208. What are the numbers? Just checking my answers, I got 8 and 12.

OpenStudy (anonymous):

x and x+4. sum of squares: 2x^2 + 8x + 16 = 208 2x^2 + 8x - 192 = 0 2(x^2 + 4x - 96) = 0 2(x - 8)(x + 12) = 0 roots at x = 8 and x = -12. throw away the negative root and we get the two integers are 8 and 12.